OpenStudy (anonymous):

2 part question for f(x)=sqroot x and g(x)=x-3 1. find the function of (f/g)(x) 2. state the domain using interval notation

OpenStudy (michele_laino):

by definition, we can write: \[\left( {\frac{f}{g}} \right)\left( x \right) = \frac{{f\left( x \right)}}{{g\left( x \right)}}\]

OpenStudy (jeff666):

f(x) = sqrt(x) g(x) = [sqrt(1-x)]^3 = (1-x)*sqrt(1-x).

OpenStudy (anonymous):

so 

OpenStudy (michele_laino):

that's right! @mlees1969

OpenStudy (jeff666):

yes there aree many ways to do it but i prefer the easy way

OpenStudy (michele_laino):

now, the radical does exist if the subsequent condition holds: \[x \geqslant 0\] and we can't divide by zero, so we get this n: \[x \ne 3\] so what can you conclude?

OpenStudy (anonymous):

and would the interval notation be